OpenStudy (anonymous):

A ground to ground projectile has a range of 200 m and time of flight of 10s and ratio of its maximum to minimum velocity during its flight is r then r is b/w (a) 1.5-2 (b) 2.5 - 3

OpenStudy (anonymous):

b

OpenStudy (anonymous):

@ajprincess

OpenStudy (anonymous):

@maheshmeghwal9 @Ishaan94

OpenStudy (anonymous):

@phi

OpenStudy (maheshmeghwal9):

ratio of max velocity with wt?

OpenStudy (anonymous):

max-min

OpenStudy (maheshmeghwal9):

oh !

OpenStudy (maheshmeghwal9):

\[\frac{u^2 \sin 2 \theta }{g}=200. \]\[\frac{2u \sin \theta }{g}=10.\] Get u & theta. then Maximum velocity is "u". Minimum velocity is "u cos theta". {at the highest point.} So ratio will be \[\frac{u}{u \cos \theta}.\]\[\implies \frac{1}{\cos \theta}.\]

OpenStudy (maheshmeghwal9):

I think it must be like this but nt sure:) but u can try my way:)

OpenStudy (maheshmeghwal9):

wt answer do u get after solving ?

OpenStudy (anonymous):

I got r= 0.05...just divide eqn 1 by eqn 2 in maheshmeghwal9's post...I don't know why the answer isn't in either one of the opns..

OpenStudy (anonymous):

The answer should be (b) 2.5 - 3

OpenStudy (maheshmeghwal9):

Main Problem is that wt is max velocity & min. velocity can anyone help us Plz?????

OpenStudy (maheshmeghwal9):

I m nt sure with my values of max velocity & min. velocity :/

OpenStudy (anonymous):

You're right about max and min velocity... As if you resolve into the components, then horiz comp remains the same alsways.., and vertical comp of velocity keeps decreasing uptil max height is reached..so till then, the velocity will decrease and after that, because the parabola is symmetric, the velocity will increase downwards and end up with the same initial value( talking about magnitudes here)..

OpenStudy (maheshmeghwal9):

thanx @aish_premrenu :D but why we aren't getting answer:/

OpenStudy (maheshmeghwal9):

can anyone help us ?

OpenStudy (experimentx):

man ... i tried of your projectiles. find the vertical component of velocity ...

OpenStudy (maheshmeghwal9):

i think that is tired*

OpenStudy (experimentx):